OpenStudy (anonymous):

Solve for h S =2(pi)rh + 2(pi)r^2

OpenStudy (anonymous):

2piR(h+r) = S S/2piR=h+r \[(s/2 \pi r) -r = h\]

OpenStudy (anonymous):

thanks brah, but i have a question what happened to the ^2

OpenStudy (anonymous):

It was factored in the first step along with the pi and the r.